Graph the line y=1/2x-4
Answer by ewatrrr(24785)


Hi
y=1/2x-4


y = (1/2)x - 4 | m = 1/2 and the y-intercept is Pt(0,-4)
Plotting that point and using the slope to find at least one more Pt.
Note: as y increases/decreases by 1, x increases/decreases by 2
Pt(2,-3) at point on the line